# It Cannot Be Right If It Was Written by AI: On Lawyers’ Preferences of Documents Perceived as Authored by an LLM vs a Human

### v1

@@ 4.2 Correctness
Formal correctness is the most frequently mentioned theme. In general, this includes correct dates and times, correct terms and conditions, identification of parties and related documents (contracts) or other external references. Formal correctness also reflects comments on the completeness or incompleteness of the information included, not only those required by the legal regulation - thus, compared to the theme of legal requirements, formal correctness is more general and broader, e.g. in a positive sense, ”the document contains all relevant information (although it does not explicitly state that the debtor is the lessee, but the content makes this clear)” or similarly in a negative sense, such as ”the document does not identify who is the creditor and who is the debtor with respect to the lease agreement” or ”it does not refer directly to the article of the agreement, where the obligation to pay rent is contained” or the more complex formulation ”it seems to me that the document correctly and consistently identifies the relevant statutory provision, the parties to the contract, and sufficiently defines the reason and the amount of the acknowledged debt”.
Formal correctness was also assessed in terms of inaccuracies or inconsistencies, e.g. ”again, I did not notice any inaccuracies, no inconsistencies either”, but also in negative terms such as ”the document is inconsistent - it does not refer to the parties and legal provisions wherever necessary” or ”the wording ’the amount will be paid in accordance with the payment terms set out in the contract’ is perhaps unnecessarily imprecise, it could be written in what specific way it will be paid”.
@@
As mentioned above, the expected legal consequences are often mentioned in the context of compliance with legal requirements, as in practice these two situations are logically linked. As with legal requirements, this is only mentioned positively, because if documents objectively meet the legal requirements, there must be expected legal consequences. Included in this theme are comments on whether the legal consequences are apparent from the document, as well as comments on the possible use of the documents in practice. Comments such as ”[…] and it is also obvious what legal effects it is intended to produce”, or ”it could probably be used in practice after minor modifications” or ”[…] although the intended legal consequences are deducible from the legal action” and ”applicable in practice”.
The final topic concerning the legal correctness of the document is that of wording and conventions. This is defined simply as the document follows the usual conventions in wording and use of legal language (in general). This is a theme that is mentioned mostly negatively, with a few positive exceptions. It is only mentioned positively when a human-crafted document is labeled as Verbose, and it is described as ”more expertly worded than an automated document” and ”the wording is concise, spare but consistent in a way that leaves no room for doubt”. In a negative sense, the specific comments regarding wording and convention are as follows: ”the phrase ’to repay’ may evoke payment in instalments - rather, to reimburse” or more general criticisms such as ”the form is skeletal” or ”the sentence ’Above: This is a debt of 25 000 CZK’ does not appear professional” and ”the sentences are oddly worded to make it clear what the document is saying” or ”the last paragraph appears ’artificial’ but contains identical information to a human-written document”.
@@
From Figure 7, it can be seen that the AI-generated document has a lower ratio of positive mentions and much higher ratio of negative ones. This is in line with the outcome of the quantitative evaluation presented above. Furthermore, there is a clear trend in Figure 7 where the negative mentions of themes for documents labeled as AI-generated outnumber the negative mentions of the themes for documents labeled as human-crafted. Conversely, the positive mentions are less common for the documents when labeled as AI-generated. Interestingly, when the brief document was labeled as generated by AI there were numerous positive mentions of the legal requirements and the expected legal consequences themes. We found that the themes are often mentioned in conjunction with a negative comment. Although, the evaluator had reservations about a document, in their opinion it met the legal requirements and would have caused the expected legal consequences.
Figure 7: Thematic analysis: Correctness, individual themes
4.3 Language
@@
Figure 8 presents the overall evaluation of the documents in terms of their language quality, focusing on whether a document was labeled as human-crafted or generated by AI. It appears that the participants clearly preferred a document when it was presented as human-crafted over a document labeled as AI-generated. Specifically, the mean evaluation of the two documents when labeled as human-crafted was 4.55 as compared to 3.97 when the same documents were marked as AI-generated. The side-by-side comparison presents a clear message when the human-crafted designation yielded 43 preferences over the mere 6 of the documents perceived as generated by AI (26 neither). This difference is statistically significant by Fisher exact test ($p<10^{-5}$).
@@
Figure 8: The figure summarizes the participants’ preferences between the documents when labeled as AI-generated versus human-crafted in terms of their Language. The top two charts show the distribution of scores awarded to the documents carrying the “AI“ (green) or “human“ (blue) labels (1–worst; 5–best). The bottom chart presents the results of the side-by side comparison, showing how many times each of the labels was preferred (if any). Overall, a clear preference for the documents labeled as human-crafted over those labeled as AI-generated can be observed.
@@
Figure 9 provides insight into the interaction between the AI/human-authored designation and the individual documents. Both documents appear to be clearly preferred by the participants when labeled as human-crafted over AI-generated. On average, the brief document was rated at 4.44 when labeled as coming from a human and at 4.11 when labeled as coming from an AI. For the verbose document, the difference was even more pronounced (4.64 human versus 3.85 AI). In terms of the side-by-side comparisons, the verbose document (labeled as AI-generated) was preferred only 1 time whereas the brief document marked as authored by a human was deemed of higher language quality 27 times (11 neither). When the brief document was marked as produced by AI it was preferred in 5 instances while the verbose document was preferred 16 times (i.e., 15 neither).
@@
Figure 9: The figure summarizes the participants’ preferences between the documents when labeled as AI-generated versus human-crafted in terms of their Language when the document is taken into account (Brief and Verbose). The top two charts show the distribution of scores awarded to the documents carrying the “AI“ (green) or “human“ (blue) labels (1–worst; 5–best). The bottom charts present the results of the side-by side comparisons, showing how many times each of the labels was preferred (if any) per document.
@@
Figure 10 focuses on the interaction between the participants’ background (i.e., law student versus lawyer) and the AI/human designation of a document. The law students clearly favored a document marked as human-authored (4.71 labeled as human versus 4.05 for AI). Further, the law students preferred the document marked as coming from a human in 25 instances compared to only 2 preferences for a document presented as generated by AI (15 neither). As in the case of correctness criterion, the lawyers were less susceptible to the AI/human designation while still preferring a document when perceived as coming from a human (4.33 human vs 3.88 AI). They expressed preference for a human-authored document in 18 cases compared to 4 instances when a document marked as AI-generated was preferred (11 neither). Note that the result is not statistically significant by Fisher exact test ($p=0.388$).
Figure 10: Correctness per background
@@
To further understand the participants’ preferences, we again conducted a thematic analysis of their explanations of the submitted document scores in terms of their language qualities. The methodology is described in Subsection 3.3. We identified four prominent themes in the explanations:
@@
  • Grammar – comments on whether the whole document is grammatically correct or whether it contains grammatical errors, such as a repetition of the reflexive pronoun in the brief document.
  • Stylistics – comments on the use of correct terminology and comments on the professionalism of the terms used.
@@
Each theme could further be mentioned in a positive (e.g., “I found no grammatical errors,” or “stylistically the document is fine”) or a negative (“I do not find the sentences easy to understand” or “the header should appear above the first paragraph of the document”) sense.
@@
Grammar is a theme containing exclusively comments on grammatical correctness or grammatical errors. In this respect, it is important to note that the Brief document intentionally contained a single grammatical error - the double reversible pronoun ’se’, specificity of the Czech language. Thus, in the case of a negative comment on grammar, the comments exclusively pointed out this error, in wording such as ”repeating ’se’ in the sentence (last paragraph)” or ”double use of the word ’se’”.
In other cases, the grammar was commented on positively as being faultless, such as ”I did not notice any grammatical or stylistic errors” or ”the document contains basically no grammatical or linguistic errors” or ”the spelling of the text is absolutely faultless’ and ’there are no spelling errors in the text”.
@@
The theme of stylistics includes references to terminology and, most often, to the expertise of legal language. Such comments are, for example, in the positive sense of ”no colloquial expressions are used (on the contrary, professional legal language is used)” or ”the text is stylistically correct and professional legal language is used” or ”stylistically, the text is perfectly fine, in terms of professionalism, the text is also fine”.
From a negative point of view, comments are, for example, ”stylistically, I would modify the document so that the line does not end with a conjunction, some sentences repeat words that could be replaced by others” or ”stylistically, there are some distracting elements, such as the excessive use of parentheses or repetitive words” and ”I consider it stylistically superfluous to use the demonstrative pronoun ’of this Agreement’ if the Agreement is previously defined as an abbreviation”.
@@
The document structure theme includes comments on missing structural parts of documents (such as headers), logical continuity of individual parts, sentence length and flow of text, document formatting and clarity (e.g. indentation). The theme of structure in a positive sense appears much more frequently in the Verbose document than in the Brief one. Specifically, these comments include, for example, ”sentences were not unnecessarily long and the complexity of the sentences did not exceed that of a standard legal text in difficulty” or ”this document is clearer and in some aspects more detailed” or ”the structure of the text is clear” and ”the structure and formatting further facilitate understanding of the text”.
From a negative point of view, the structure is mentioned, for example, as ”the sequence of the text, where only in the 3rd paragraph it is mentioned that the debtor acknowledges his debt, seems illogical to me” or ”there is a lack of a header for better orientation as regards the pages of the acknowledgement in the text, some sentences are too long” or ”for the sake of clarity, I would have placed the heading at the head of the document, not in the text”. It is also commented as ”slightly distracting to me, however, is the way the text is indented (first line of all paragraphs)” and ”not quite standard formatting”.
@@
The topic of clarity is more often mentioned in a positive sense for a document labeled as human-crafted, while AI-generated is more often rated negatively in terms of clarity. In general, clarity comments refer to the accuracy and comprehensibility of the expression. These include e.g. ”the document is completely and totally comprehensible” or ”the text is much more readable” or simple ”it is comprehensible”. In negative terms, the comments relate to the excessive complexity of the text to e.g. ”I find the text of the document unnecessarily complex” or ”the document could be edited to make it simpler and easier to understand” or ”it is, for a ’normal reading’, a complex text”.
@@
The ratio of positive versus negative mentions confirms the results of the quantitative analysis. The detailed results of the thematic analysis are shown in Figure 11. There, we can see the high prevalence of negative mentions for the documents when labeled as AI-generated (both verbose and brief), across all the themes. When the documents are labeled as human-crafted positive mentions prevail. We also find a notable difference in the ratings for the brief document. When labeled as AI-generated it contains a significantly higher number of negative comments on language quality in each category.
Figure 11: Thematic analysis: Language, individual themes
4.4 Beliefs in the Possibility of Automatic Generation
@@
After evaluating the documents, the participants were asked to answer an open-ended question on the possibility of the fully automated generation of the documents similar to those they have just seen. The overwhelming majority of 93% of the participants answered that they believed this is going to be possible. Only 5% remained uncertain, and merely 2% believed this to be impossible. We further explored the reasons that led the evaluators to believe in the possibility of future automation. Responses to this question were thematically coded following the methodology described in Subsections 3.3. The codes were then collated into five higher-level themes as follows:
@@
  • Documents Indistinguishable – the evaluators claimed that they found human-crafted and AI-generated documents practically indistinguishable and thus full automation was possible.
@@
  • Expected Legal Consequences – the evaluators argued that if, despite some reservations, the document is capable of producing the expected legal consequences, then it is very likely that full automation of such documents is possible.
@@
  • Similar to Forms – the evaluators argued that it is already possible to mechanize some legal actions by using form legal documents (e.g. contracts) and this is essentially more advanced work with such forms.
@@
  • Sufficient Content – the evaluators commented that they found the content of the document sufficient and so did not see a problem with full automation.
@@
  • Lower Cost – the evaluators commented on the cost/benefit by comparing the time and money saved by automatic generation even at the cost of the document not being perfect.
@@
Figure 12: Thematic analysis: Reasons for possibility of future automation
A general conclusion in the context of the negative evaluation of both the language and the correctness of the document, which was labeled AI-generated, is that even though the participants evaluated such a document more negatively, they believe that it would still hold up in practice and therefore consider the automation of legal documents to be very realistic. Despite the evaluators’ reservations about the linguistic and legal quality of a document labeled as AI-generated, it can be assumed from their answers that the document is in their opinion sound and can be fully automated.

### v2

@@ 4.2 Correctness
Formal correctness is the most frequently mentioned theme. It includes correct dates and times, terms and conditions, identification of parties and related documents (contracts) or other external references. Formal correctness also contains comments on the completeness or incompleteness of the information included, not only those required by the legal regulation - thus, compared to the theme of legal requirements, formal correctness is more general and broader. In a positive sense, e.g.:
  • ”The document contains all relevant information (although it does not explicitly state that the debtor is the tenant, but the content makes this clear).”
  • ”It seems to me that the document correctly and consistently identifies the relevant statutory provision, the parties to the contract, and sufficiently defines the reason and the amount of the acknowledged debt.”
On the other hand, negative comments appeared to focus on a lack of detail, e.g.,
  • ”The document does not identify the creditor and the debtor concerning the lease agreement.”
  • ”It does not refer directly to the article of the agreement, where the obligation to pay rent is contained.”
Formal correctness was also assessed regarding inaccuracies or inconsistencies. Positive mentions included a lack of any inaccuracies and inconsistencies. On the other hand, some participants reported negatively emphasizing inconsistencies in referring to parties and legal provisions or lacking details about the debt payment method.
@@
As mentioned above, the expected legal consequences are often mentioned in the context of compliance with legal requirements, as in practice, these two situations are logically linked. As with legal requirements, the theme is only mentioned positively because if documents objectively meet the legal requirements, there must be expected legal consequences. The theme includes comments on whether the legal consequences are apparent from the document and the possible use of the documents in practice. Comments such as
  • ”[…] and it is also obvious what legal effects it intends to cause.”
  • ”It could probably be used in practice after minor modifications.”
  • ”[It is] applicable in practice.”
The final topic concerning the correctness of the document is that of wording and conventions. This is defined simply as the document following the usual conventions in wording and use of legal language (in general). The theme is mentioned mostly negatively, with a few positive exceptions. It is only discussed positively when a Verbose document is labeled as human-crafted. In that case, it is described as
  • ”More expertly worded than an automated document.”
  • ”The wording is concise, spare, but consistent in a way that leaves no room for doubt.”
In a negative sense, comments focused on inappropriate form, lack of professionalism, or poor choice of words. Examples include,
  • ”The form is skeletal.”
  • ”The sentence ’This is a debt of 25 000 CZK’ does not appear professional.”
  • ”The sentences are oddly worded to be absolutely certain about the document’s meaning.”
  • ”The last paragraph appears ’artificial’ but contains identical information to a human-written document.”
@@
Figure 7 shows that the AI-generated document has a lower ratio of positive mentions and a much higher ratio of negative ones. Note that the results of thematic analysis are not tested for statistical significance, but intended to demonstrate patterns and trends. The thematic analysis aligns with the outcome of the quantitative evaluation presented above. Furthermore, there is a clear trend in Figure 7 where the negative mentions of themes for documents labeled as AI-generated outnumber the negative mentions of the themes for documents labeled as human-crafted. Conversely, positive mentions are less common for documents when labeled as AI-generated. Interestingly, when the Brief document was labeled as generated by AI, there were numerous positive mentions of the legal requirements and the expected legal consequences themes. We found these themes often mentioned in conjunction with a negative comment. Although the evaluator had reservations about a document, in their opinion, it met the legal requirements and would have caused the expected legal consequences.
Figure 7: Thematic analysis: Frequency of positive and negative mentions for each individual theme related to Correctness
4.3 Language Quality
@@
Figure 8 presents the overall evaluation of the documents in terms of their language quality, focusing on whether a document was labeled as human-crafted or AI-generated. The participants preferred a document when it was presented as human-crafted over a document labeled as AI-generated. Specifically, the mean evaluation of the two documents when labeled as human-crafted was 4.55 as compared to 3.97 when the same documents were marked as AI-generated. The side-by-side comparison presents a clear message when the human-crafted designation yielded 43 preferences over the mere 6 of the documents labeled as AI-generated (26 neither). This difference is statistically significant by Fisher exact test ($p<10^{-4}$ using Bonferroni correction for 4 comparisons).
@@
Figure 8: The figure summarizes the participants’ preferences between the documents when labeled as AI-generated versus human-crafted in terms of their Language Quality. The top two charts show the distribution of scores awarded to the documents carrying the “AI“ (green) or “human“ (blue) labels (1–worst; 5–best). The bottom chart presents the results of the side-by side comparison, showing how many times each of the labels was preferred (if any). Overall, a clear preference for the documents labeled as human-crafted over those labeled as AI-generated can be observed.
@@
Figure 9 provides insight into the interaction between the AI/human-authored designation and the individual documents. Both documents appear to be preferred by the participants when labeled as human-crafted over AI-generated. On average, the Brief document was rated at 4.44 when labeled as coming from a human and at 4.11 when labeled as coming from an AI. For the Verbose document, the difference was even more pronounced (4.64 human versus 3.85 AI). In terms of the side-by-side comparisons, the Verbose document (labeled as AI-generated) was preferred only 1 time. In contrast, the Brief document marked as human-crafted was deemed of higher language quality 27 times (11 neither). When the Brief document was marked as produced by AI, it was preferred in 5 instances, while the Verbose document was preferred 16 times (i.e., 15 neither).
@@
Figure 9: The figure summarizes the participants’ preferences between the documents when labeled as AI-generated versus human-crafted in terms of their Language Quality when the document is taken into account (Brief and Verbose). The top two charts show the distribution of scores awarded to the documents carrying the “AI“ (green) or “human“ (blue) labels (1–worst; 5–best). The bottom charts present the results of the side-by side comparisons, showing how many times each of the labels was preferred (if any) per document.
@@
Figure 10 focuses on the interaction between the participants’ background (i.e., law student versus lawyer) and the AI/human designation of a document. Note that the result is not statistically significant by the Fisher exact test ($p=0.388$). The law students appeared to favor a document marked as human-crafted (4.71 labeled as human versus 4.05 for AI). Further, the law students preferred the document marked as coming from a human in 25 instances compared to only 2 preferences for a document presented as generated by AI (15 neither). As in the case of the correctness criterion, the lawyers appeared to have been less susceptible to the AI/human designation while still preferring a document when perceived as coming from a human (4.33 human vs 3.88 AI). They preferred a human-authored document in 18 cases compared to 4 instances when a document marked as AI-generated was preferred (11 neither).
Figure 10: The figure summarizes the participants’ preferences between the documents when labeled as AI-generated versus human-crafted in terms of their Language Quality when the participants’ background is taken into account (a law student and a lawyer). The top two charts show the distribution of scores awarded to the documents carrying the “AI“ (green) or “human“ (blue) labels (1–worst; 5–best). The bottom charts present the results of the side-by-side comparisons, showing how many times each of the labels was preferred (if any) per participants’ background. It appears the AI-generated label has larger effect on the law students as compared to the lawyers.
@@
To further understand the participants’ preferences, we again conducted a thematic analysis of their explanations of the submitted document scores regarding their language qualities. The methodology is described in Subsection 3.3. We identified four prominent themes in the explanations:
@@
  • Grammar – comments on whether the whole document is grammatically correct or contains grammatical errors, such as the repeating reflexive pronoun in the Brief document.
  • Stylistics – comments on using the correct professional terminology.
@@
Each theme could be mentioned positively, e.g.,
  • “I found no grammatical errors.”
  • “Stylistically speaking, the document is fine.”
It could also be mentioned negatively, such as
  • “I find the sentences difficult to understand.”
  • “The header should appear above the first paragraph of the document.”
@@
Grammar is a theme containing exclusively comments on grammatical correctness or errors. In this respect, it is worth noting the Brief document intentionally contained a single grammatical error - the double reversible pronoun ’se’, specificity of the Czech language. Thus, in the case of a negative comment on grammar, the comments exclusively pointed out this error in wording.
In other cases, the grammar was commented on positively as being flawless, such as
  • ”I did not notice any grammatical or stylistic errors”
  • ”The document contains no grammatical or linguistic errors.”
  • ”The spelling of the text is flawless.”
  • ”there are no spelling errors in the text.”
@@
The theme of stylistics includes references to terminology and, most often, to the professionalism of legal language. Positive comments include
  • ”No colloquial expressions are used (on the contrary, professional legal language is used).”
  • ”The text is stylistically correct, and professional legal language is used.”
  • ”Stylistically, the text is excellent, in terms of professionalism, the text is also fine.”
Negative comments on the theme appeared as well, such as
  • ”Stylistically, I would modify the document so that the line does not end with a conjunction. Some sentences repeat words that could be replaced by alternatives.”
  • ”Stylistically, there are some distracting elements, such as the excessive use of parentheses or repetition of words.”
  • ”I consider it stylistically superfluous to use the demonstrative pronoun ’of this Agreement’ if the Agreement is previously defined as an abbreviation.”
@@
The document structure theme includes comments on missing structural parts of documents (such as headers), logical continuity of individual parts, sentence length and flow of text, document formatting and clarity (e.g. indentation). In a positive sense, the theme of structure appears much more frequently in the Verbose document than in the Brief one. Specifically, these comments include,
  • ”Sentences were not unnecessarily long, and the complexity of the sentences did not exceed that of a standard legal text in difficulty.”
  • ”The structure of the text is clear.”
  • ”The structure and formatting further facilitate understanding of the text.”
Negatively, the theme structure is mentioned, for example, as
  • ”The sequence of the text, where only the 3rd paragraph mentions the debtor acknowledging the debt, seems illogical to me.”
  • ”There is a lack of a header for better orientation as regards the parties of the acknowledgement in the text. Some sentences are too long.”
  • ”For the sake of clarity, I would have placed the header ahead of the document, not in the text.”
  • ”The way the text is indented is slightly distracting to me.”
@@
The theme of clarity is more often mentioned positively for a document labeled as human-crafted, while AI-generated is more often mentioned negatively. In general, clarity comments refer to the accuracy and comprehensibility of the expression. In negative terms, the comments relate to the excessive complexity of the text and the possibility of making it simple through editorial intervention.
@@
The ratio of positive versus negative mentions confirms the results of the quantitative analysis. Note that the results of thematic analysis are not tested for statistical significance, but intended to demonstrate patterns and trends. The detailed results of the thematic analysis are shown in Figure 11. There, we can see the high prevalence of negative mentions for the documents when labeled as AI-generated (both Verbose and Brief) across all the themes. Once the documents are labeled as human-crafted, positive mentions prevail. We also find a notable difference in the ratings for the Brief document. When labeled as AI-generated, more negative comments on language quality in each category appeared.
Figure 11: Thematic analysis: Frequency of positive and negative mentions for each individual theme related to Language Quality,
4.4 Possibility of Automated Generation
@@
After evaluating the documents, the participants were asked to answer an open-ended question about the possibility of a fully automated generation of documents similar to those they had just seen. The overwhelming majority, 93%, answered that they believed this would be possible. Only 5% remained uncertain, and merely 2% believed this to be impossible.
We further explored the reasons that led the participants to believe in the possibility of future automation. Responses to this question were thematically coded following the methodology described in Subsections 3.3. The codes were then collated into five higher-level themes as follows:
@@
  • Documents Indistinguishable – the evaluators claimed that they found human-crafted and AI-generated documents practically indistinguishable, and thus full automation was possible.
@@
  • Expected Legal Consequences – the evaluators argued that if, despite some reservations, the document can produce the expected legal consequences, then it is very likely that full automation of such documents is possible.
@@
  • Similarity to Forms – the evaluators argued that it is already possible to mechanize some legal actions by using forms (e.g. simple contracts), and using LLMs is akin to more technologically advanced forms.
@@
  • Sufficient Content – the evaluators commented that they found the content of the document sufficient and, as a result, did not see a problem with full automation.
@@
  • Lower Cost – the evaluators commented on the cost/benefit by comparing the time and money saved by automatic generation, even at the cost of the document not being perfect.
@@
Figure 12: Thematic analysis: Frequency of reasons for possibility of future automation
A general conclusion in the context of the negative evaluation of both the language and the correctness of the document, which was labeled AI-generated, is that even though the participants evaluated such a document more negatively, they believe it would still hold up in practice. Therefore, they consider the automation of legal documents to be very realistic. Despite the evaluators’ reservations about the linguistic and legal quality of a document labeled as AI-generated, it can be assumed from their answers that the document is, in their opinion, sound and can be fully automated.